# Some insights into depth estimators for location and scatter in the multivariate setting

### v1

@@ 5 Appendix
$\inf_{\Sigma\in\mathcal{F}\left(M\right),P\in\mathcal{P}_{\varepsilon}\left(F_{0}\right)}P\left(\left\|\hat{\Sigma}-\Sigma\right\|_{op}^{2}\leq C^{\ast}\left(\max\left\{\frac{p}{n},B_{E}^{2}(\varepsilon)\right\}+\frac{\log\left(1/\delta\right)}{n}\right)\right)\geq\alpha.$
@@ 5.1 Maximum bias of the deepest scatter estimator
We know that $\hat{\Gamma}\left(P_{0}\right)$ is Fisher consistent up to a constant, that is $\hat{\Gamma}\left(P_{0}\right)=\sqrt{\beta}\Sigma.$ Without loss of generality to study the breakdown point and maximum bias, we can assume that $\Sigma=I.$ If $F$ is a centrosymmetric distribution in $\mathbb{R}^{p}$ and $X\sim F\,$with then $\mathbf{w}^{t}X\sim\mathbf{z}^{t}X$ for all $\mathbf{w,z\in\mathcal{S}}^{p-1}.$ Let us take a symmetric positive matrix $\Gamma,$with eigenvalues $l_{1}\geq l_{2}\geq\dots$ $\geq l_{p}>0$ and eigenvectors $\mathbf{v}_{1},\dots,\mathbf{v}_{p}$ respectively
$\Gamma=\sum_{j=1}^{p}l_{j}\mathbf{v}_{j}\mathbf{v}_{j}^{t}.$
Observe that
$P_{0}\left(\left|\mathbf{u}^{t}X\right|^{2}\leq\mathbf{u}^{t}\Gamma\mathbf{u}\right)$ $=$ $P_{0}\left(\left|\mathbf{u}^{t}X\right|^{2}\leq\mathbf{u}^{t}\Gamma\mathbf{u}\right)=P_{0}\left(-\sqrt{\mathbf{u}^{t}\Gamma\mathbf{u}}\leq\mathbf{u}^{t}X\leq\sqrt{\mathbf{u}^{t}\Gamma\mathbf{u}}\right)$
$P_{0}\left(\left|\mathbf{u}^{t}X\right|^{2}\geq\mathbf{u}^{t}\Gamma\mathbf{u}\right)$ $=$ $P_{0}\left(\mathbf{u}^{t}X\leq-\sqrt{\mathbf{u}^{t}\Gamma\mathbf{u}}\right)+P\left(\mathbf{u}^{t}X\geq\sqrt{\mathbf{u}^{t}\Gamma\mathbf{u}}\right),$
Let $g$ be the function,
$\dot{g}$ ::: $\emph{S}^{p-1}\rightarrow\left[0,1\right],$
$g\left(\mathbf{u}\right)$ $=$ $P_{0}\left(-\sqrt{\mathbf{u}^{t}\Gamma\mathbf{u}}\leq\mathbf{u}^{t}X\leq\sqrt{\mathbf{u}^{t}\Gamma\mathbf{u}}\right).$
Thus, we have the following result.
Lemma 11
$\ D\left(\Gamma,P_{0}\right)=\min_{\mathbf{u\in}\emph{S}^{p-1}}\left(g\left(\mathbf{u}\right),1-g\left(\mathbf{u}\right)\right)=\min\left(g\left(\mathbf{v}_{p}\right),1-g\left(\mathbf{v}_{1}\right)\right)$
Proof. Given $g\left(\mathbf{u}\right)=\Phi\left(\sqrt{\mathbf{u}^{t}\Gamma\mathbf{u}}\right)-\Phi\left(-\sqrt{\mathbf{u}^{t}\Gamma\mathbf{u}}\right),$ the Lagrangian is
@@
[…]
Therefore, we get that
@@
and the minimum is given by $m=\min\left(g\left(\mathbf{v}_{p}\right),1-g\left(\mathbf{v}_{1}\right)\right).$
Corollary 12
The deepest estimator is a multiple of the identity and its depth is $1/2$
Proof. Since $g\left(\mathbf{v}_{p}\right)\leq g\left(\mathbf{v}_{1}\right)$ and are increasing functions of $\sqrt{\mathbf{v}^{t}\Gamma\mathbf{v}}$, then $g\left(\mathbf{v}_{p}\right)=1-g\left(\mathbf{v}_{1}\right)=0.5$ which says. $l_{p}^{1/2}=l_{1}^{1/2}=\Phi^{-1}\left(3/4\right).$ Then we $\Gamma=\left[\Phi^{-1}\left(3/4\right)\right]^{2}I_{p}$ ($I_{p}$ the identity matrix $p\times p).$
Corollary 13
Let $\left\{P_{\varepsilon,n}\right\}_{n=1}^{\infty}$ be such that $P_{\varepsilon,n}=\left(1-\varepsilon\right)P_{0}+\varepsilon P_{n}$ such that the associated depth estimators $\left\{\Gamma_{n}\right\}_{n=1}^{\infty}$, $\Gamma_{n}=\sum_{j=1}^{p}l_{j}^{(n)}\mathbf{v}_{j}^{(n)}\left[\mathbf{v}_{j}^{(n)}\right]$ have either their largest eigenvalue $l_{1}^{(n)}\rightarrow\infty$ or their smallest eigenvalue $l_{p}^{(n)}\rightarrow 0.$ Then $\varepsilon>1/3.$
Proof. Put
$a_{\varepsilon,n}\left(\mathbf{w}\right)$ $=$ $\left(1-\varepsilon\right)P_{0}\left(\frac{\mathbf{w}^{t}\Gamma_{n}^{-1/2}}{\left\|\mathbf{w}^{t}\Gamma_{n}^{-1/2}\right\|}XX^{t}\frac{\Gamma_{n}^{-1/2}\mathbf{w}}{\left\|\mathbf{w}^{t}\Gamma_{n}^{-1/2}\right\|}\leq\frac{1}{\sum_{j=1}^{p}l_{j}^{{}^{(n)}-1}\left(\mathbf{w}^{t}\mathbf{e}_{j}\right)^{2}}\right)+$
$\varepsilon P_{n}\left(\frac{\mathbf{w}^{t}\Gamma_{n}^{-1/2}}{\left\|\mathbf{w}^{t}\Gamma_{n}^{-1/2}\right\|}XX^{t}\frac{\Gamma_{n}^{-1/2}\mathbf{w}_{r}}{\left\|\mathbf{w}^{t}\Gamma_{n}^{-1/2}\right\|}\leq\frac{1}{\sum_{j=1}^{p}l_{j}^{{}^{(n)}-1}\left(\mathbf{w}^{t}\mathbf{e}_{j}\right)^{2}}\right)$
$b_{\varepsilon,n}\left(\mathbf{w}\right)$ $=$ $\left(1-\varepsilon\right)P_{0}\left(\frac{\mathbf{w}_{r}^{t}\Gamma_{n}^{-1/2}}{\left\|\mathbf{w}^{t}\Gamma_{n}^{-1/2}\right\|}XX^{t}\frac{\Gamma_{n}^{-1/2}\mathbf{w}_{r}}{\left\|\mathbf{w}^{t}\Gamma_{n}^{-1/2}\right\|}\geq\frac{1}{\sum_{j=1}^{p}l_{j}^{{}^{(n)}-1}\left(\mathbf{w}^{t}\mathbf{e}_{j}\right)^{2}}\right)+$
$\varepsilon P_{n}\left(\frac{\mathbf{w}_{r}^{t}\Gamma_{n}^{-1/2}}{\left\|\mathbf{w}_{r}^{t}\Gamma_{n}^{-1/2}\right\|}XX^{t}\frac{\Gamma_{n}^{-1/2}\mathbf{w}_{r}}{\left\|\mathbf{w}^{t}\Gamma_{n}^{-1/2}\right\|}\geq\frac{1}{\sum_{j=1}^{p}l_{j}^{{}^{(n)}-1}\left(\mathbf{w}^{t}\mathbf{e}_{j}\right)^{2}}\right)$
Let $\mathbf{w}_{n\text{ }}$ be such that $D\left(\Gamma_{n},P_{n}\right)=\min\left(a_{\varepsilon,n}\left(\mathbf{w}_{n}\right),b_{\varepsilon,n}\left(\mathbf{w}_{n}\right)\right).$
@@
i) If $l_{1}^{(n)}\rightarrow\infty,$ since $\min\left(a_{\varepsilon,n}\left(\mathbf{v}_{1}^{(n)}\right),b_{\varepsilon,n}\left(\mathbf{v}_{1}^{(n)}\right)\right)\geq\min\left(a_{\varepsilon,n}\left(\mathbf{w}_{n}\right),b_{\varepsilon,n}\left(\mathbf{w}_{n}\right)\right)$ then, either $b_{\varepsilon,n}\left(\mathbf{v}_{1}^{(n)}\right)\geq a_{\varepsilon,n}\left(\mathbf{v}_{1}^{(n)}\right)\geq D\left(\Gamma_{n},P_{n}\right)$ or $a_{\varepsilon,n}\left(\mathbf{v}_{1}^{(n)}\right)\geq b_{\varepsilon,n}\left(\mathbf{v}_{1}^{(n)}\right)\geq D\left(\Gamma_{n},P_{n}\right)$ which yields that
$\left(1-\varepsilon\right)P_{0}\left(\frac{\mathbf{w}_{n}^{t}\Gamma_{n}^{-1/2}}{\left\|\mathbf{w}_{n}^{t}\Gamma_{n}^{-1/2}\right\|}XX^{t}\frac{\Gamma_{n}^{-1/2}\mathbf{w}_{r}}{\left\|\mathbf{w}_{n}^{t}\Gamma_{n}^{-1/2}\right\|}\geq l_{1}^{(n)}\right)$
$+\varepsilon P_{n}\left(\frac{\mathbf{w}_{n}^{t}\Gamma_{n}^{-1/2}}{\left\|\mathbf{w}^{t}\Gamma_{n}^{-1/2}\right\|}XX^{t}\frac{\Gamma_{n}^{-1/2}\mathbf{w}_{n}}{\left\|\mathbf{w}^{t}\Gamma_{n}^{-1/2}\right\|}\geq l_{1}^{(n)}\right)$
$\geq$ $D\left(\Gamma_{n},P_{n}\right)$
Observe that $P_{0}\left(\frac{\mathbf{w}_{n}^{t}\Gamma_{n}^{-1/2}}{\left\|\mathbf{w}^{t}\Gamma_{n}^{-1/2}\right\|}XX^{t}\frac{\Gamma_{n}^{-1/2}\mathbf{w}_{r}}{\left\|\mathbf{w}^{t}\Gamma_{n}^{-1/2}\right\|}\geq l_{1}^{(n)}\right)\rightarrow 0$ and we get

### v2

@@ 7 Appendix
$\inf_{\Sigma\in\mathcal{F}\left(M\right),P\in\mathcal{P}_{\varepsilon}\left(P_{0}\right)}P\left(\left\|\hat{\Sigma}-\Sigma\right\|_{op}^{2}\leq C^{\ast}\left(\max\left\{\frac{p}{n},B_{E}^{2}(\varepsilon)\right\}+\frac{\log\left(1/\delta\right)}{n}\right)\right)\geq\alpha.$
[…]
Proof of Lemma 3.4. Since we have regression through the origin model (a zero intercept model) and the covariables have an ellipsoidal distribution and we are dealing with affine equivariant estimators we can assume that $B=0,$ $\Sigma=I$ and $\sigma=1$, Then, $P_{B},$ the joint distribution of $\left(\mathbf{X},\mathbf{Y}\right)\in\mathbb{R}^{p}\times\mathbb{R}^{m},$ is given by
$\left(\begin{array}[]{c}\mathbf{Y}\\ \mathbf{X}\end{array}\right)\sim N_{p+m}\left(\left(\begin{array}[]{c}0_{m\times 1}\\ 0_{p\times 1}\end{array}\right),\left(\begin{array}[]{cc}I_{m\times m}&0_{m\times p}\\ 0_{p\times m}&I_{p\times p}\end{array}\right)\right)$
Then, if $A,U\in\mathbb{R}^{p\times m},$ it holds that
$\left(\begin{tabular}[]{l}$\mathbf{Y}-A^{t}\mathbf{X}$\\ $U^{t}\mathbf{X}$\end{tabular}\right)\sim N_{2m}\left(\left(\begin{array}[]{c}0_{m\times 1}\\ 0_{m\times 1}\end{array}\right),\left(\begin{array}[]{cc}I_{m\times m}+B^{t}B&-B^{t}U\\ -U^{t}B&U^{t}U\end{array}\right)\right).$
Given $D\left(A,P_{B}\right)=\inf_{U\in\mathbb{R}^{p\times m}}P_{B}\left((U^{t}X,Y-A^{t}X\right)\geq 0),$ from the proof of Theorem 4.1, p.1167 of Gao (2020), we have the following inequality
$D\left(\hat{A},P_{B}\right)\geq\frac{1}{2}-\frac{\varepsilon}{1-\varepsilon}-C_{1}\left(\sqrt{\frac{mp+p}{n}}+\sqrt{\frac{\log\left(1/\delta\right)}{2n}}\right),$ (29)
with $C_{1}$ an absolute constant (independent of $n,p,m,\varepsilon)$). Then,
$P_{B}\left((U^{t}X,Y-A^{t}X\right)\geq 0)$ $=$
$=$ $E_{X}\left[P\left(\left.X^{t}UA^{t}X\leq X^{t}UY\right|X\right)\right]\text{ \ }(\left.X^{t}UY\right|X\sim N\left(0,\left\|X^{t}U\right\|^{2}\right)$
$=$ $E\left[1-\Phi\left(\frac{X^{t}UA^{t}X}{\left\|X^{t}U\right\|}\right)\right]=1-E\Phi\left(\left(\frac{U^{t}X}{\left\|U^{t}X\right\|},A^{t}X\right)\right).$
Therefore,
$\inf_{U\in\mathbb{R}^{p\times m}}P_{B}\left((U^{t}X,Y-A^{t}X\right)$ $\geq$ $0)=1-\sup_{U\in\mathbb{R}^{p\times m}}E\Phi\left(\left(\frac{U^{t}X}{\left\|U^{t}X\right\|},A^{t}X\right)\right)$
$=$ $1-E\Phi\left(\left\|A^{t}X\right\|\right).$
Take the spectral decomposition of the non-negative matrix $A^{t}\in\mathbb{R}^{p\times p}, AA$$A^{t}=PDP^{t}$ $P\in\mathbb{R}^{p\times p}$ orthonormal and $D=diag\left(\lambda_{1},...,\lambda_{p}\right)$ with $0\leq\lambda_{0}\leq\lambda_{1}\leq\dots\leq\lambda_{p}$, then $Z=PX\sim N_{p}\left(0,I\right).$ Therefore,
$E\Phi\left(\sqrt{X^{t}AA^{t}X}\right)=E\Phi\left(\sqrt{Z^{t}DZ}\right)=E\Phi\left(t\sqrt{Z^{t}\frac{D}{tr\left(D\right)}Z}\right)=E\Phi\left(t\sqrt{\sum_{i=1}^{p}\theta_{i}Z_{i}^{2}}\right),$
with $t=\sqrt{tr\left(D\right)},$ $\left\{\theta_{i}\right\}_{i=1}^{p}\subset\left[0,1\right]$ and $\sum_{i=1}^{p}\theta_{i}=1$. The square root is a concave function, and we can obtain that
$\sqrt{\sum_{i=1}^{p}\theta_{i}Z_{i}^{2}}\geq\sum_{i=1}^{p}\theta_{i}\left|Z_{i}\right|.$
$\Phi$ is also a concave function on $0,\infty)$, then take the functions $h:[0,\infty)\times\left[0,1\right]^{p}\rightarrow\left[0.5,1\right]$ and $g:[0,\infty)\rightarrow\left[0.5,1\right]$ as $h\left(t,\theta_{1},\dots,\theta_{p}\right)=\Phi\left(t\sqrt{\sum_{i=1}^{p}\theta_{i}Z_{i}^{2}}\right)$ and $g\left(t\right)=h\left(t,1,0,...,0\right),$
$\Phi\left(t\sqrt{\sum_{i=1}^{p}\theta_{i}Z_{i}^{2}}\right)$ $\geq$ $\sum_{i=1}^{p}\theta_{i}\Phi\left(t\left|Z_{i}\right|\right)$
$E\Phi\left(t\sqrt{\sum_{i=1}^{p}\theta_{i}Z_{i}^{2}}\right)$ $\geq$ $\sum_{i=1}^{p}\theta_{i}E\Phi\left(t\left|Z_{i}\right|\right)=E\Phi\left(t\left|Z_{1}\right|\right)=g\left(t\right),$
since $E\Phi\left(t\left|Z_{1}\right|\right)=\dots=E\Phi\left(t\left|Z_{p}\right|\right).$ Then
$\inf_{U\in\mathbb{R}^{p\times m}}P_{B}\left((U^{t}X,Y-A^{t}X\right)\geq 0)=1-E\Phi\left(t\sqrt{\sum_{i=1}^{p}\theta_{i}Z_{i}^{2}}\right)\leq 1-E\Phi\left(t\sqrt{Z_{1}^{2}}\right)=1-g\left(t\right).$
Then, from (29) we can say,
$D\left(\hat{A},P_{B}\right)$ $=$ $1-g\left(\hat{t}\right)\geq\frac{1-3\varepsilon}{2\left(1-\varepsilon\right)}-C_{1}\left(\sqrt{\frac{mp+p}{n}}+\sqrt{\frac{\log\left(1/\delta\right)}{2n}}\right)$
[…]
Since $g$ is increasing, we get that
$\hat{t} \displaystyle\leq$ $g^{-1}\left(\frac{1+\varepsilon}{2\left(1-\varepsilon\right)}+C_{1}\left(\sqrt{\frac{pm+p}{n}}+\sqrt{\frac{\log\left(1/\delta\right)}{2n}}\right)\right)$
$=$ $g^{-1}\left(\frac{1+\varepsilon}{2\left(1-\varepsilon\right)}+C_{1}\left(\sqrt{\frac{pm+p}{n}}+\sqrt{\frac{\log\left(1/\delta\right)}{2n}}\right)\right)-g^{-1}\left(\frac{1+\varepsilon}{2\left(1-\varepsilon\right)}\right)+g^{-1}\left(\frac{1+\varepsilon}{2\left(1-\varepsilon\right)}\right)$
$=$ $C_{1}\left(g^{-1}\right)^{\prime}\left(\xi\right)\left(\sqrt{\frac{pm+p}{n}}+\sqrt{\frac{\log\left(1/\delta\right)}{2n}}\right)+g^{-1}\left(\frac{1+\varepsilon}{2\left(1-\varepsilon\right)}\right),\text{ }$
$\xi$ $\in$ $\left(\frac{1+\varepsilon}{2\left(1-\varepsilon\right)},\frac{1+\varepsilon}{2\left(1-\varepsilon\right)}+C_{1}\left(\sqrt{\frac{pm+p}{n}}+\sqrt{\frac{\log\left(1/\delta\right)}{2n}}\right)\right).$
Thus,
$\hat{t}\leq C\left[\left(\frac{pm}{n}\vee g^{-1}\left(\frac{1+\varepsilon}{2\left(1-\varepsilon\right)}\right)\right)+\sqrt{\frac{\log\left(1/\delta\right)}{2n}}\right]$
[…]
Remark 7.1
Observe that $g^{-1}\left(\frac{1+\varepsilon}{2\left(1-\varepsilon\right)}\right)$ does not depend either on the number of dependent variables $m$ or the number of independent variables $p.$
Proof of Lemma 3.5. Since we are dealing with affine equivariant estimators we can assume that $\boldsymbol{\beta}=0,$ $\Sigma=I$ and $\sigma=1$. With this background we get that
$\left(\begin{array}[]{c}y\\ \mathbf{X}\end{array}\right)\sim N_{p+1}\left(\left(\begin{array}[]{c}0\\ \mathbf{0}_{p\times 1}\end{array}\right),\left(\begin{array}[]{cc}1&\mathbf{0}_{p\times 1}^{t}\\ \mathbf{0}_{p\times 1}&I_{p\times p}\end{array}\right)\right)$
Then, it holds that
$\left(\mathbf{\lambda}^{t}\mathbf{X},y-\mathbf{\alpha}^{t}\mathbf{X}\right)^{t}\sim N_{2}\left(\left(\begin{array}[]{c}0\\ 0\end{array}\right),\left(\begin{array}[]{cc}1&-\mathbf{\alpha}^{t}\mathbf{\lambda}\\ -\mathbf{\alpha}^{t}\mathbf{\lambda}&1+\left\|\mathbf{\alpha}\right\|^{2}\end{array}\right)\right)$
for a unitary vector $\mathbf{\lambda}.$ Since
$P_{B}\left(\left(\mathbf{\lambda}^{t}\mathbf{X}\right)(y-\mathbf{X}^{t}\mathbf{\alpha})\right)\geq 0)=1-E\Phi\left(\frac{\mathbf{\lambda}^{t}\mathbf{XX}^{t}\mathbf{\alpha}}{\sigma\left|\mathbf{\lambda}^{t}\mathbf{X}\right|}\right).$
Since $h\left(-\rho\right)=1-h\left(\rho\right)$ and $\dot{h}:\left[-1,1\right]\rightarrow\left[0,1\right]$ is increasing, we have that, for $b=\left\|\alpha\right\|$
$P_{B}\left(\left(\mathbf{\lambda}^{t}\mathbf{X}\right)(y-\mathbf{X}^{t}\mathbf{\alpha})\right)$ $\geq$ $0)=1-P_{B}\left(\left(\mathbf{\lambda}^{t}\mathbf{X}\right)(y-\mathbf{X}^{t}\mathbf{\alpha})\right)<0)$
$=$ $1-\left[1-h\left(-\frac{\mathbf{\alpha}^{t}\mathbf{\lambda}}{\sqrt{1+b^{2}}}\right)\right]$
$=$ $h\left(-\frac{\mathbf{\alpha}^{t}\mathbf{\lambda}}{\sqrt{1+b^{2}}}\right)=1-h\left(\frac{\mathbf{\alpha}^{t}\mathbf{\lambda}}{\sqrt{1+b^{2}}}\right)$
This entails that,
$\inf_{\left\|\mathbf{\lambda}\right\|=1}P_{B}\left(\left(\mathbf{\lambda}^{t}\mathbf{X}\right)(y-\mathbf{X}^{t}\mathbf{\alpha})\right)$ $\geq$ $0)=1-\sup_{\left\|\lambda\right\|=1}E\Phi\left(\frac{\mathbf{\lambda}^{t}\mathbf{XX}^{t}\mathbf{\alpha}}{\left|\mathbf{\lambda}^{t}\mathbf{X}\right|}\right)$
$=$ $1-E\Phi\left(\left\|\mathbf{\alpha}\right\|\frac{\left|\mathbf{\alpha}^{t}\mathbf{X}\right|}{\left\|\mathbf{\alpha}\right\|}\right)=1-E\Phi\left(\left\|\mathbf{\alpha}\right\|\left|Z\right|\right)=1-g\left(\left\|\mathbf{\alpha}\right\|\right)$
$=$ $1-\sup_{\left\|\lambda\right\|=1}h\left(\frac{\mathbf{\alpha}^{t}\mathbf{\lambda}}{\sqrt{1+b^{2}}}\right)=1-h\left(\frac{b}{\sqrt{1+b^{2}}}\right)$
Then,
$h\left(\frac{b}{\sqrt{1+b^{2}}}\right)=g\left(b\right)=\frac{1+\varepsilon}{2\left(1-\varepsilon\right)},$
[…]
@@ 7.1 Maximum bias of the deepest scatter estimator
Proof of Lemma 4.1. Given $g\left(\mathbf{u}\right)=\Phi\left(\sqrt{\mathbf{u}^{t}\Gamma\mathbf{u}}\right)-\Phi\left(-\sqrt{\mathbf{u}^{t}\Gamma\mathbf{u}}\right),$ the Lagrangian is given by
@@
$0$ $\mathbf{=}$ $\frac{\partial h\left(\mathbf{u,}\lambda\right)}{\partial\lambda}=\mathbf{u}^{t}\mathbf{u-}1.$
Therefore, we obtain that
@@
[…]
Proof of Corollary 4.2 Since $g(\mathbf{v}_{p})\leq g(\mathbf{v}_{1})$ and $g$ is an increasing function of $\sqrt{\mathbf{v}^{\top}\Gamma\mathbf{v}}$, $\min\bigl(g(\mathbf{v}_{p}),\,1-g(\mathbf{v}_{1})\bigr)$ is maximized when $g(\mathbf{v}_{p})=1-g(\mathbf{v}_{1})=0.5$, which implies $\ell_{p}^{1/2}=\ell_{1}^{1/2}=\Phi^{-1}(3/4)$. Hence, $\hat{\Gamma}=\left[\Phi^{-1}(3/4)\right]^{2}I_{p},$ and the corresponding maximum depth equals to $1/2$.
Proof of Corollary 4.3 For every vector $\mathbf{w}\in\mathbb{R}^{p}$, define
$Z_{n}(\mathbf{w})=\frac{\mathbf{w}^{t}\Gamma_{n}^{-1/2}\mathbf{X}}{\left\|\mathbf{w}^{t}\Gamma_{n}^{-1/2}\right\|}\in\mathbb{R}.$
Let
$c_{n}(\mathbf{w})=\frac{1}{\displaystyle\sum_{j=1}^{p}l_{j}^{(n)-1}\left(\mathbf{w}^{t}\mathbf{e}_{j}\right)^{2}}.$
and consider the probabilities
$a_{\varepsilon,n}(\mathbf{w})$ $=$ $(1-\varepsilon)\,P_{0}\!\left(Z_{n}(\mathbf{w})^{2}\leq c_{n}(\mathbf{w})\right)+\varepsilon\,P_{n}\!\left(Z_{n}(\mathbf{w}_{r})^{2}\leq c_{n}(\mathbf{w})\right),$
$b_{\varepsilon,n}(\mathbf{w})$ $=$ $(1-\varepsilon)\,P_{0}\!\left(Z_{n}(\mathbf{w}_{r})^{2}\geq c_{n}(\mathbf{w})\right)+\varepsilon\,P_{n}\!\left(Z_{n}(\mathbf{w}_{r})^{2}\geq c_{n}(\mathbf{w})\right).$
[…]
Case i) $l_{1}^{(n)}\rightarrow\infty$.
By definition of the depth, $D(\Gamma_{n},P_{n})=\min\{a_{\varepsilon,n}(\mathbf{w}_{n}),b_{\varepsilon,n}(\mathbf{w}_{n})\}\leq\min\{a_{\varepsilon,n}(\mathbf{w}),b_{\varepsilon,n}(\mathbf{w})\}$ for every vector $\mathbf{w}$. Since
$D(\Gamma_{n},P_{n})=\min\{a_{\varepsilon,n}(\mathbf{w}_{n}),b_{\varepsilon,n}(\mathbf{w}_{n})\}\leq\min\{a_{\varepsilon,n}(\mathbf{v}_{1}^{(n)}),b_{\varepsilon,n}(\mathbf{v}_{1}^{(n)})\},$
it holds that $a_{\varepsilon,n}(\mathbf{v}_{1}^{(n)})\geq D(\Gamma_{n},P_{n})\quad b_{\varepsilon,n}(\mathbf{v}_{1}^{(n)})\geq D(\Gamma_{n},P_{n})$. By considering that $\{\mathbf{v}_{j}^{(n)}\}_{j=1}^{p}$ is an orthonormal basis, we have that
$\frac{1}{\sum_{j=1}^{p}l_{j}^{(n)-1}\bigl(\mathbf{v}_{1}^{(n)T}\mathbf{e}_{j}\bigr)^{2}}=l_{1}^{(n)}.$
and
$b_{\varepsilon,n}\!\left(\mathbf{v}_{1}^{(n)}\right)$ $=$ $(1-\varepsilon)\,P_{0}\!\left(Z_{n}\!\left(\mathbf{v}_{1}^{(n)}\right)^{2}\geq l_{1}^{(n)}\right)+\varepsilon\,P_{n}\!\left(Z_{n}\!\left(\mathbf{v}_{1}^{(n)}\right)^{2}\geq l_{1}^{(n)}\right)$
$\geq$ $D(\Gamma_{n},P_{n}).$
@@
Since $P_{0}\left(Z_{n}\!\left(\mathbf{v}_{1}^{(n)}\right)^{2}\geq l_{1}^{(n)}\right)\rightarrow 0$ when $l_{1}^{(n)}\rightarrow\infty$, we may conclude that $\varepsilon\geq$ $\lim_{n\rightarrow\infty}D\left(\Gamma_{n},P_{n}\right)\geq$ $\lim_{n\rightarrow\infty}D\left(I,P_{n}\right)\geq$ $\left(1-\varepsilon\right)D\left(I,P_{0}\right)=\left(1-\varepsilon\right)\frac{1}{2}$which says that $\varepsilon\geq 1/3.$
Case ii) $l_{p}^{(n)}\rightarrow 0$. With a similar analysis to the case i) we can conclude that
$b_{\varepsilon,n}\left(\mathbf{v}_{p}^{(n)}\right)\geq D\left(\Gamma_{n},P_{n}\right)\text{ and }a_{\varepsilon,n}\left(\mathbf{v}_{p}^{(n)}\right)\geq D\left(\Gamma_{n},P_{n}\right).$
Since
$a_{\varepsilon,n}\left(\mathbf{v}_{p}^{(n)}\right)=(1-\varepsilon)\,P_{0}\!\left(Z_{n}\!\left(\mathbf{v}_{p}^{(n)}\right)^{2}\leq l_{p}^{(n)}\right)+\varepsilon\,P_{n}\!\left(Z_{n}\!\left(\mathbf{v}_{p}^{(n)}\right)^{2}\leq l_{p}^{(n)}\right)\geq D\!\left(\Gamma_{n},P_{n}\right).$
and considering that $P_{0}\left(Z_{n}\!\left(\mathbf{v}_{p}^{(n)}\right)^{2}\leq l_{p}^{(n)}\right)\rightarrow 0$ when $l_{p}^{(n)}\rightarrow 0$, we get